### v8

# Latency Adjustable Transformer Encoder for Language Understanding

Abstract
Adjusting the latency, power, and accuracy of natural language understanding models is a desirable objective of an efficient architecture. This paper proposes an efficient Transformer architecture that adjusts the inference computational cost adaptively with a desired inference latency speedup. In fine-tuning phase, the proposed method detects less important hidden sequence elements (word-vectors) and eliminates them in each encoder layer using a proposed Attention Context Contribution (ACC) metric. After the fine-tuning phase, with the novel offline-tuning property, the inference latency of the model can be adjusted in a wide range of inference speedup selections without any further training. The proposed method is applied to the BERTbase, GPT-2 and Flan-T5 models for evaluation. Extensive experiments show that most of the word-vectors in higher Transformer layers have less contribution to the subsequent layers; hence, they can be eliminated to improve the inference latency. Experimental results on extensive sentiment analysis, classification, text generation tasks and regression benchmarks like GLUE showed that the method is effective in various datasets with minimal impact on the input’s global context. The method was also evaluated under the instruction tuning paradigm, and its performance was measured using different types of prompting. The proposed method mathematically and experimentally improves the inference latency of BERTbase and GPT-2 by up to 4.8 and 3.72 times with less than 0.75% accuracy drop and passable perplexity on average. The suggested approach posits that in Large Language Models (LLMs), although the complete network is necessary for training, it can be truncated during the fine-tuning phase.
Index Terms:
Word-vector elimination profile, attention context contribution (ACC), offline-tuning, latency adjustable.
I Introduction
Recently, Transformer [1] based architectures achieved remarkable success in various Natural Language Processing (NLP) tasks. However, the training difficulty, inference latency, and small size of available datasets are the main concerns with using these models in real applications [2]. Fine-tuning a pre-trained language model on downstream tasks mitigates some of these concerns, including training effort and dataset size. GPT [3], BERT [4], and XLNet [5] are examples of recent state-of-the-art pre-trained language models. The number of parameters in these models is more than hundreds of millions. Hence, the computational complexity bears billions of Floating Point Operations (FLOPs) for processing a single sample. In some cases, the inference phase computational effort is beyond the ability of resource-limited systems such as edge devices. Therefore, baseline Transformer models cause unacceptable latency and energy usage limitations in such devices.
One of the most effective pre-trained language models is Bidirectional Encoder Representations from Transformers (BERT) [4], which has multi-layer Transformer-based architecture. BERTbase consists of an embedding layer and 12 stacked encoder layers. Each encoder contains an attention layer with 12 heads and multiple fully connected layers. There are about 110M parameters in BERTbase model, the pre-training of which has taken 79 hours on 64 Tesla V100 GPUs [2].
In the realm of Deep Neural Networks (DNNs) optimization, diverse methods have been proposed to enhance speed and reduce power consumption. This is crucial for applications such as extending battery life and enabling powerful tasks on small devices. The approaches can be categorized into pruning, removing unnecessary parts; quantization, simplifying numerical usage; Knowledge Distillation, simplifying based on a complex model; parameter-sharing, using the same information across components; tensor decomposition, breaking down data structures; and sub-quadratic complexity transformers [6].
This study introduces an innovative attention layer designed to enhance the efficiency of inference stages in Transformer-based models with an adjustable approach. The suggested method involves a sorting and elimination process for hidden sequence elements (word-vectors), resulting in a reduction of effective word-vectors in each layer and a decrease in the number of Floating Point Operations (FLOPs). The primary contributions of this research can be summarized as follows:
  • The novel structure and tuning policy are proposed to speed up the inference times with negligible accuracy degradation.
  • A new Attention Context Contribution (ACC) metric is proposed to observe the context contribution in each attention layer. It is shown that the context contribution is decreased at the last encoder layers significantly.
  • An accurate analytical inference speedup estimation is presented for the proposed method that helps a system designer determine the hyper-parameter selection at the fine-tuning phase.
  • Offline-tuning property that adjusts the inference latency after the fine-tuning phase. This property controls the trade-off between latency and accuracy in a wide range of inference speedup selections without requiring additional training.
  • The proposed approach can be applied to diverse encoder and decoder Transformer-based architectures and preserves the input’s global context.
II Related Works
Pruning-based methodologies have been extensively explored in the existing body of literature, with categorizations based on distinct pruning approaches. In weight pruning methods, the model’s weights are systematically pruned according to a predetermined threshold [7, 8, 9, 10]. Head and channel pruning methods extend their scope by entirely excising a channel or head of the network [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Layer and block pruning strategies [22, 23, 24], involve the targeted removal of entire layers or blocks within the network. Token and word-vector pruning methods, adopting a more fine-grained approach, focus on the selective removal of specific tokens or word-vectors within the model [25, 26, 27]. This classification clarifies the various approaches taken in the pruning paradigm, each playing a unique role in achieving the overall goal of optimizing networks.
II-A Weight Pruning
In [28], the magnitude weight pruning method is applied to BERT and shows that the model can be pruned once during pre-training rather than separately for each task without affecting performance. Reweighted Proximal Pruning (RPP) has introduced a pruning method specifically designed for a large-scale language representation model that is much better than typical iterative pruning. This approach yields a sparse version of BERT [29]. Transformer mixed-length vector pruning (TF-MVP) examines the sparsity patterns in pruned Transformers and introduces the TF-MVP architecture using the mixed-length vector pruning (MVP) technique [30]. While TF-MVP achieves a pruning ratio of approximately 70%, there is a notable decline in accuracy. In [31], the model weights are encouraged to be zero with the L0 regularization term.
II-B Head/Channel Pruning
A parameter reduction method was proposed by identifying the most critical heads in each encoder layer using layer-wise relevance propagation and pruning the redundant heads [32]. It showed that the earlier encoder layer’s heads are much more critical than the last. Moreover, in [33], it is shown that around 17% of attention heads can be removed at the test time without significantly impacting inference performance. In [34], a rapid post-training pruning framework tailored for Transformers is introduced. This framework involves three key fine-tuning stages: mask search, mask rearrangement, and mask tuning. Notably, the method obviates the necessity for any retraining, presenting an efficient approach to achieve model compression in the Transformer architecture.
II-C Layer/Block Pruning
A progressive layer-dropping at the pre-training phase of BERT architecture offers up to 2.5 times speedup at the inference phase with more knowledge transferability and better generalization on downstream tasks [35]. FastBERT is proposed with adaptive inference time in which the model uses a sufficient number of encoder layers based on the inference certainty of a sample [36]. It dynamically adjusts the number of active layers to reduce the overall computational effort. ELBERT proposed the Confidence-Window Based Early Exit mechanism, improving the ALBERT inference speed [37]. In [38], it is experimentally shown that using all layers of a pre-trained model in downstream tasks is unnecessary. Studying various layer-dropping strategies shows that many model weights are not necessary. LayerDrop introduced a type of structured dropout that imparts regularization benefits during the training process and enables effective pruning at inference time, resulting in the removal of approximately half of the Transformer layers [39].
II-D Token/Word-Vector Pruning
PoWER-BERT proposes a word-vector elimination method that eliminates redundant word-vectors[25]. A supplement learnable layer is added between each encoder self-attention and feed-forward layer, and an additional loss function term is used to train these supplement parameters. PoWER-BERT training consists of 3 steps “fine-tuning, elimination configuration search, and re-training” [25]. Funnel-Transformer is inspired by U-net from computer vision [40]. It improves the Transformer architecture by inserting a pooling layer in each self-attention layer [26]. Funnel-Transformer compresses the sequence of hidden states to a shorter one and reduces the computation cost. Learned Token Pruning (LTP) adaptively removes unimportant tokens as an input sequence passes through transformer layers [27]. LTP is applied to RoBERTa [41], which results in up to 2.0 times throughput improvement with less than 1% accuracy drop.
In STTABT [42], the impact of token pruning on later layers’ attentions and on final predictions is studied. The study shows that the proposed attention back-tracking allows the model to better retain the full model’s performance even at high sparsity rates. Conversely, since in the self-attention architecture all input elements contribute to the formation of each attention output, it can be stated that in the word-vector pruning procedure, no independent concept is discarded. In the subsequent layers, every concept retains the ability to expand and influence the outputs and model predictions. Additionally, STTABT prunes tokens based on a lightweight attention approximation network trained with knowledge distillation, which adds extra training and inference overhead.
In [43], a learnable mechanism is employed to determine which uninformative tokens can be dropped from the context at any point during the generation process. In [44], the model leverages an asymmetric attention mechanism to iteratively distill inputs into a tight latent bottleneck, allowing it to scale to handle very large inputs.
II-E Non-Pruning Methods
ALBERT is a lighter version of BERT model [45]. It uses cross-layer weight sharing and a modified embedding layer structure, making the embedding size independent of the encoder’s hidden size. Moreover, the new Sentence Order Prediction (SOP) task is utilized in the pre-training phase instead of the Next Sentence Prediction task. ALBERT reports state-of-the-art results with its parameter reduction and modified pre-training policy. ALBERT is nine times smaller and 1.2 times faster than BERTbase, with less than 2.5% on average accuracy drop. ALBERT is an example of the parameter-sharing technique.
DistilBERT [46], TinyBERT [47] and MobileBERT [48] use the knowledge distillation technique during the pre-training and offer up to 5.5 times faster inference on downstream tasks than BERTbase. In [49] a Self-Distillation (SD) mechanism obtains high-accuracy models directly without going through an assistive model. The main concerns with these methods are the pre-training effort of the model from scratch, extra training steps, and significant accuracy degradation.
In [50], the authors propose sufficient conditions under which they prove that a sparse attention model can universally approximate any sequence-to-sequence function. They show that sparse Transformers with only $\mathcal{O}_{n}$ connections per attention layer can approximate the same function class as the dense model with $n^{2}$ connections. Consequently, in recent years, many works [51, 52, 53, 54, 55, 56, 57] have proposed methods that aim to reduce the complexity of the attention mechanism from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$, where $N$ is the sequence length.
Low-Rank Adaptation (LoRA) involves the freezing of weights in the pre-trained model, and during the fine-tuning phase, it introduces trainable rank decomposition matrices into each layer of the Transformer architecture [58]. LoRA enhances the efficiency of memory usage during the training phase by up to 3 times, but it does not confer any advantages during the inference phase. QLoRA improves over LoRA by quantizing the transformer model to 4-bit precision and using paged optimizers to handle memory spikes [59]. LQ-LoRA uses an iterative algorithm to decompose each pre-trained matrix into a high-precision low-rank component and a memory-efficient quantized component [60].
The previous works are trying to reduce the model size and inference latency. These works show that significant part computations in the models appear unnecessary, and models with less computational complexity can be obtained from the original ones. The main drawbacks of these methods are pre-training from scratch, multi-phase fine-tuning, and significant inference accuracy drop. In these methods, the inference speedup is set in the fine-tuning phase, and the achieved speedup is fixed in the inference phase. The proposed method tries to mitigate the previous drawbacks. It presents the powerful offline-tuning mechanism that controls the inference speedup of the model after the fine-tuning phase.
The impact of the speedup methods on the model stability and the input’s global context understanding is an important concern that must be studied in methods evaluations. Backpack learns multiple non-contextual sense vectors for each word in a vocabulary and represents a word in a sequence as a context-dependent, non-negative linear combination of sense vectors in this sequence [61]. The Backpack Model shows that each word in the input sequence can interpreted as a sense vector that encodes different aspects of a word which is very sensitive. A good practice for assessing the performance of a method is evaluating it in a more complex task, like text generation tasks that inspect the model proficiency. Therefore the proposed method is applied to the GPT-2 model beside the BERTbase model and evaluated on several text generation tasks.
III Proposed Method
This Section presents details of the proposed method and the analysis and reasoning behind the method. Section III-A presents the preliminary background of the BERTbase architecture, and Section III-B analyzes the FLOPs in the BERTbase model. The proposed encoder layer is based on an insight into the amount of the contribution of word-vectors at different layers, which is considered in Section III-C. In Section III-D, the proposed architecture is investigated in detail. In Section III-E, the proposed method is extended to decoder architecture and applied to the GPT-2 model.
III-A Background
Fig. 1 shows BERTbase model architecture and output classifier. The input sentence goes through the tokenizer layer and embedding layer. Then the context is fed into 12 cascaded encoder layers to be processed. The pooler and classifier layers are added in the last layer to perform proper output for the desired task. In the figure, $T$ and $H$ are the numbers of input word-vectors and the hidden state vector size, respectively. Each encoder consists of a Bidirectional Multi-Head Self-Attention layer and a Feed-Forward layer with several tensor multiplications. The residual branches and normalization layers increase the training stability.
Figure 1: BERTbase [4] architecture with output layers
III-B BERT FLOPs Analysis
Table I shows the estimated FLOPs of BERTbase at different layers. The amount of FLOPs is estimated based on the number of input tokens ($T$), the number of encoder layers ($L$), the hidden state size ($H$), and the number of output labels ($N$).
TABLE I: BERTbase FLOPs Study
Layer Sub-layer FLOPs Estimationa FLOPs Estimation FLOPs Shareb Latency Sharec
Embedding Word-embeddings 0
Position-embeddings 0 $7TH$ 0.003% 0.2%
Layer-normalization $7TH$
Encoder (12 layers) Attention-self $6LTH^{2}$ + $2HTL^{2}$ $6LTH^{2}$ + $2HTL^{2}$ 25.1%
Attention-output $2LTH^{2}$ 99.8%
Intermediate $8LTH^{2}$ $18LTH^{2}$ 74.9%
Output $8LTH^{2}$
Classifier Pooler $2H^{2}$ $2H^{2}$ 0.0013% 0.036%
Output $2HN$
a $T$ = number of input tokens, $L$ = number of encoder layers, $H$ = hidden state size and $N$ = number of output labels
b With $T$ = 512, $L$ = 12, $H$ = 768 and $N$ = 2
c IMDB [62] dataset average relative inference time latency.
[…]
The inference latency and computational effort are directly related to the number of FLOPs. As a result, reducing the number of FLOPs at the encoder layers reduces the total number of FLOPs and inference latency.
III-C Attention Context Contribution (ACC) metric
The context contribution of word-vectors at each self-attention layer is studied in this Section to find more important word-vectors at each encoder layer. For a given encoder layer, the Attention Context Contribution (ACC) metric measures the contribution of word-vectors in the layer’s output. This metric is derived from a Score Vector (SV) to find more important word-vectors in each layer. For a given layer, the SV indicates the contribution of each input word-vector to the output of the attention layer.
Fig. 2 illustrates the calculation of the Score Vector based on the attention probability matrix. Fig. 2 (left) shows the attention probability matrix of BERTbase model. The rows of this matrix are normalized to one. By averaging the attention probability matrix over the heads, the aggregated probability matrix is resulted, Fig. 2 (center). The SV results from the sum of this matrix along the unnormalized dimension as shown in Fig. 2 (right). The SV elements with higher score values correspond to the word-vectors that contribute more to the layer output.
Figure 2: Obtaining Score Vector (SV) from Attention Probability Matrix in a given BERT [4] encoder layer
Fig. 3 illustrates BERTbase SV results on the IMDB sentiment analysis task[62] for an input sample. As shown in earlier encoder layers, the SV values approximately have uniform distributions, and in the last encoder layers, the distribution of SV values changes to approximately $\delta$ distribution. The distribution of SV values in the last layers indicates that the concentration of scores is high on a few word-vectors and others have less contribution to the self-attention layer output.
Figure 3: BERTbase[4] Score Vector (SV) results on the IMDB sentiment analysis task[62]
The SV elements are positive and with an average of one. So, the median of SV is bounded. Suppose the median of the SV is lower than its average. In that case, more than half of the elements are lower than the average value, and only a few are higher than the average value. It suggests that most of the word-vectors do not bear enough contribution to the layer output, and only a few word-vectors influence the final result of the layer. The median of the Score Vector is interpreted as the proposed Attention Context Contribution metric.
The ACC metric calculation is explained in Algorithm. 1. The algorithm inputs are BERT encoder word-vectors and the model parameters. The algorithm output is the ACC metric vector that suggests the contribution of the word-vectors in each encoder self-attention layer. For encoder layer $l$, the first step is to prepare the pre-computed self-attention probability matrix; $M_{Att}^{l}$. Then the average of this matrix over its heads is computed; $M_{H}^{l}$. In the next step, the resulting matrix’s summation over its unnormalized dimension is computed; $M_{S}^{l}$. This matrix is the Score Vector. Attention Context Contribution of the layer ($E_{ACC}^{l}$) is the median of the Score Vector.
Algorithm 1 Attention Context Contribution (ACC) metric
0:  Encoder inputs $\in R^{LTH}$ fine-tuned BERT {$T$ = number of input word-vectors, $L$ = number of encoder layers, $H$ = hidden state size}
0:  ACC vector: $E_{ACC}\in R^{L}$ Initialisation:
1:  $E_{ACC}=[\emptyset]$ LOOP Process
2:  for $encoder\ l$ in $Encoders$ do
3:     $M_{Att}^{l}\leftarrow$ Self-Attention Probability Matrix $\in R^{HTT}$
4:     $M_{H}^{l}\leftarrow$ Mean of $M_{Att}^{l}$ over heads $\in R^{TT}$
5:     $M_{S}^{l}\leftarrow$ Sum of $M_{H}^{l}$ over unnormalized dimension $\in R^{T}$
6:     $E_{ACC}^{l}\leftarrow$ Median of $M_{S}^{l}$ (Score Vector) $\in R$
7:     Append $E_{ACC}^{l}$ to $E_{ACC}$
8:  end for
9:  return  $E_{ACC}$
Fig. 4 shows the ACC metric for BERTbase encoder layers on IMDB [62] dataset and several General Language Understanding Evaluation (GLUE) [63] benchmark tasks. The fitted curve to the ACC metric results shows that the ACC metric is reduced at later layers gradually. It indicates that the fraction of the word-vectors that contribute more in the last layers is less in those layers; hence more word-vectors can be eliminated. By eliminating less important word-vectors in each layer, the processed word-vectors are decreased gradually, and the inference time of the model is decreased.
As shown in Fig. 4, the behavior of the ACC metric strongly correlates with the intricate specifications inherent in each task. The ACC results consistently exhibit a monotonic decrease with the layer number, except in the case of the SST-2 [64] dataset, where unexpected behavior is observed. Notably, the SST-2 dataset, centered around sentiment analysis of movie reviews, shares a strong resemblance to the IMDB dataset. A comparative analysis of the ACC curves for SST-2 and IMDB reveals a notable disparity, primarily associated with the length of input sequences. Specifically, in the IMDB dataset, the majority of samples consist of around 512 tokens; conversely, in the SST-2 dataset, the majority of samples feature fewer than 50 tokens. It can be inferred that the limited number of input tokens leads to an undesired sensitivity to word-vectors, and in such cases, some word-vectors are revived through residual branches.
Figure 4: BERTbase [4] Encoder Attention Context Contribution (ACC) metric results (blue bars) and second degree fitted curves (red lines) on SST-2 [64], MRPC [65], QQP [66], QNLI [67], RTE [68], MNLI [69], STS-B [70] and IMDB [62] tasks.
III-D Proposed Architecture
According to the discussion in Section III-B, almost all latency due to the FLOPs occur in the encoder layers. In addition, according to the observation in Section III-C, the word-vector contribution in the encoder decreases at later layers. Based on these observations, this Section proposes the details of the proposed architecture.
The proposed method is applied to the BERTbase model for explanation and evaluation. The proposed architecture is the same as the original BERT with a modified multi-head self-attention layer. Fig. 5 illustrates the proposed multi-head self-attention layer. The proposed architecture adds the Sort and Eliminate layers in the middle of the original BERT multi-head self-attention layer. The Sort layer sorts the input context according to the computed layer Score Vector. The Eliminate layer eliminates a fraction of the sorted context with lower scores according to the Elimination-Rate ($\alpha_{ER}$).
Figure 5: The proposed self-attention layer
$\alpha_{ER}$ is a hyper-parameter vector that sets the fraction of word-vectors that is eliminated in each encoder layer. For layer $l$, $\alpha_{ER}^{l}$ is the word-vector Elimination Rate obtained from the pre-computed Elimination-Profile ($\alpha_{EP}$) vector and a pre-specified Speedup-Coefficient hyper-parameter ($\alpha_{SC}$). $\alpha_{SC}$ adjusts the trade-off between accuracy and speed in fine-tuning and offline-tuning phases.
$\alpha_{ER}^{l}=\alpha_{EP}^{l}\times\alpha_{SC}$ (1)
$\alpha_{EP}^{l}$ is obtained from the fitted curve to the ACC metric (red lines in Fig. 4) at layer $l$. In some tasks, the ACC values of layers are not smooth. To increase stability and prevent applying stress to the model in fine-tuning phase, in the proposed formulation, instead of ACC metric values ($E_{ACC}$), the fitted curve to the ACC metric ($P_{ACC}$) is used that presents a more smooth elimination trend. It is important to note that when an increasing behavior of the ACC value occurs in a layer, the elimination process is halted, and the remaining layers proceed using a fixed number of retained word-vectors. In (2), the $\alpha_{EP}$ lower bound is set to one to ensure a monotonic descending shape of this hyper-parameter curve. This adaptive approach guarantees the preservation of crucial information and mitigates the impact of unexpected ACC behavior on subsequent layers.
$\alpha_{EP}^{l}=\min{(1,\frac{P_{ACC}^{l}}{P_{ACC}^{l-1}})}$ (2)
The number of remaining word-vectors in each encoder layer is calculated according to the number of the layer input word-vectors and the $\alpha_{ER}$ hyper-parameter. Equation (3) presents the number of remaining word-vectors ($T_{l}$) at the output of encoder layer $l$. The minimum number of remaining word-vectors at each encoder output is limited to one.
$T_{l}=\max{(1,\lfloor\alpha_{ER}^{l}\times T_{l-1}\rfloor)}$ (3)
The effective number of processed word-vectors can be estimated using the number of word-vectors before and after the eliminate layer. Based on Table I, in each encoder layer, around 75% of FLOPs occur after the self-attention layer. Equation (4) presents the effective number of Processed word-vectors ($PW_{l}$) in layer $l$.
$PW_{l}\approx\frac{T_{l-1}+3T_{l}}{4}$ (4)
Hence,
$PW\approx T\left[\frac{1}{4}+\sum_{i=1}^{L-1}\prod_{j=1}^{i}\alpha_{EP}^{j}+\frac{3}{4}\prod_{j=1}^{L}\alpha_{EP}^{j}\right]$ (5)
Processed word-vectors in the original BERTbase is:
[…]
In these equations, ${L}$ is the number of BERT encoder layers, and ${T}$ is the number of encoder input word-vectors. Since almost all inference latency occurs in the encoder, the overall model computation effort is almost equal to that of the encoder stage. Equation (7) presents the proposed method inference time speedup to BERTbase.
$K_{speedup}\approx\frac{4L}{1+4\sum_{i=1}^{L-1}\prod_{j=1}^{i}\alpha_{EP}^{j}+3\prod_{j=1}^{L}\alpha_{EP}^{j}}$ (7)
The above equation assumes that the number of remaining word-vectors in each encoder layer is more than a single word-vector. In other words, (3) returns the second argument, which is valid for most cases. For example, if $L$ is 12 and $\alpha_{EP}^{l}$ is constant and equal to 0.8 for all layers, the method inference speed is around 3.09 times faster than BERTbase.
The proposed Sort and Elimination layers are included in the fine-tuning phase. Given that the eliminated word-vectors constitute a relatively small portion within each layer during the fine-tuning process, we can assume that the gradients effectively propagate through the remaining word vectors. Additionally, the residual paths stay functional, contributing to the updates in the initial layers of the network. This design ensures a robust flow of information throughout the fine-tuning process.
To enhance the speed of the suggested self-attention layer, the Sort and Eliminate layers can be positioned after the self-attention concatenation layer. This adjustment results in a reduction of per-head calculations, and for a given layer, the eliminated word-vectors play a role in shaping the layer’s output. Furthermore, in accordance with the suggested word-vector elimination strategy, the pruned word-vectors efficiently accelerate the inference phase, leading to an anticipated close alignment between the reduction in Floating Point Operations (FLOPs) and the speedup in inference.
III-E Extending Proposed Architecture to Decoder Models
The applicability of the proposed method can be extended to non-encoder Transformers, such as the GPT-family Large Language Models (LLMs). GPT-2 [71] represents a decoder-based transformed language model widely utilized in text generation tasks. This section outlines the considerations when implementing the proposed method on the GPT-2 model.
The foundational GPT-2 model comprises 12 decoder layers, with each decoder featuring a causal attention layer equipped with 12 heads. The total number of parameters in the GPT-2 model is approximately 117 million. A key distinction between the architectures of GPT-2 and the BERTbase model lies in the design of their attention layers. As explained in Section III-A, the attention layers of BERTbase encoders are bidirectional, involving all input word-vectors in generating each output word-vector. Conversely, the GPT-2 model, tailored for text generation tasks, employs causal attention layers, where only the preceding word-vectors contribute to the attention layer output at each time stamp.
In a causal attention layer, the attention probability matrix, represented as $M_{Att}^{l}$, adopts the form of a triangular matrix. To extend the proposed architecture outlined in Section III-D to the GPT-2 model, an additional normalization step needs to be integrated after the fifth step in Algorithm 1. This supplementary step normalizes the Score Vector $M_{S}^{l}$ based on the count of non-zero elements in the probability matrix $M_{Att}^{l}$. Equation (8) articulates this normalization process in layer $l$, wherein the $i$-th element of the Score Vector is adjusted relative to the number of non-zero elements in the $i$-th column of the probability matrix. Through this normalization step, the resulting score values become comparable.
$M_{S_{i}}^{l^{\prime}}=M_{S_{i}}^{l}\div N_{MAtt_{i}}^{l}$ (8)
IV Experiments
In this section, we implement the proposed method on both BERTbase, GPT-2 and Flan-T5 [72] models, delving into the method’s effectiveness across various natural language understanding tasks, text generation datasets and instruction tuning benchmarks. Towards the conclusion of this section, we delve into the offline-tuning property and conduct an ablation study on the proposed self-attention layer.
IV-A Datasets
The proposed method performance is assessed in encoder models using the IMDB [62] dataset, which is a sentiment analysis task, as well as the General Language Understanding Evaluation (GLUE) [63] benchmark, which consists of two single-sentence classification tasks: SST-2 [64] and CoLA [73], two multi-sentence similarity classification tasks: MRPC [65] and QQP [66], one binary question answering task: QNLI [67], three natural language inference tasks: RTE [68], MNLI-M and MNLI-MM [69], and one multi-sentence similarity regression task: STS-B [70].
To assess the performance of the method in decoder models, we conduct evaluations on several datasets. The WikiText-103 [74] language modeling dataset, comprising over 100 million tokens extracted from verified Good and Featured articles on Wikipedia, serves as a primary evaluation ground. Additionally, the PTB [75] dataset, featuring a million words from 1989 Wall Street Journal material, and the 1BW [76] dataset, established as a benchmark corpus for measuring progress in statistical language modeling, are included in our evaluation. Moreover, we assess the method on The LAMBADA [77] dataset, a collection of narrative passages characterized by the intriguing feature that human subjects can predict their last word when exposed to the entire passage. Extracted from BookCorpus, The LAMBADA dataset comprises 10,022 passages, contributing to a comprehensive evaluation of the proposed method’s performance in decoder models.
To more closely show the method’s performance to understand language, it is applied to the Flan-T5 model and evaluated on the Massive Multitask Language Understanding (MMLU) benchmark, which includes exam questions from 57 tasks such as mathematics, history, law, and medicine [78], and the Big Bench Hard (BBH) benchmark, which includes 23 challenging tasks [79].
IV-B Model Settings
[…]
For the instruction tuning evaluation, the Flan-T5 model is used. Flan-T5 is an enhanced version of T5 [80] that has been fine-tuned on a mixture of tasks. The Flan-T5 encoder self-attention layers are substituted with the proposed encoder self-attention layer, and other layers are modified to work with a dynamic number of word vectors in each encoder and decoder layer. The Flan-T5 weights are directly used in our model without any fine-tuning process. the $\alpha_{ER}$ is selected in the range [0.8, 1.0], and other evaluation hyper-parameters are the same as those of the default Flan-T5 model.
IV-C Implementation Details
The initial models utilized for this study are the pre-trained versions of BERTbase-uncased [4], GPT-2 [71], and Flan-T5 [72] from the Hugging Face [^2] https://huggingface.co/ implementations. The implementation is carried out in the PyTorch [^3] https://pytorch.org/ library, utilizing an Nvidia Tesla P100 GPU for BERTbase, an Nvidia Tesla T4 GPU for GPT-2, and an Nvidia Tesla A100 GPU for Flan-T5. The datasets are sourced from the Hugging Face Datasets library, and the GLUE results are validated through the official GLUE service [^4] https://gluebenchmark.com/. The instruction tuning assessment is done with help of the INSTRUCTEVAL [81] method. This framework ensures consistency and reliability in the assessment of the proposed method’s performance on the specified models.
IV-D Experimental Results of BERTbase
The experimental results for extensive tasks are presented in Table II. This work is 2.9 times faster than BERTbase, with 0.7% drop on the test set. Also, the model inference time is up to 4.8 times faster than BERTbase, with only 0.9% accuracy drop. These results indicate that the proposed method effectively improves Transformer-based models’ performance and resource usage with minimal accuracy degradation.
Table II shows the experimental speedup, the expected speedup, and the number of FLOPs for each task. The number of FLOPs reduction is consistent with inference speedup and expected speedup predicted using (7). The slight differences are due to the computational overheads, including the computational effort of the tokenizer and output classifier layers that are not considered in (7). The consistency between the experimental speedup, the expected speedup, and the FLOPs reduction approves the hypothesis discussed in Section III-D.
TABLE II: The Proposed Method Experimental Results on GLUE [63] benchmark and IMDB [62] dataset
Model CoLA SST-2 MRPC QQP QNLI RTE MNLI-M MNLI-MM STS-B IMDB Avg
Test Resulta BERTbase 52.7 93.5 88.7 71.2 91.0 68.9 84.7 83.8 84.3 93.7 81.25
This Work 53.4 92.6 88.0 70.1 90.2 67.9 83.7 82.7 84.2 92.7 80.55
Inference Speedup - 3.8 2.4 2.9 4.7 2.3 2.6 2.5 2.5 2.3 2.7 2.9
Expected Speedupb - 3.67 2.35 2.88 4.76 2.25 2.64 2.49 2.49 2.22 2.63 2.84
#FLOPs - 5.7B 9.1B 7.6B 4.7B 9.8B 8.5B 8.7B 8.7B 9.5B 8.6B 7.7B
Input Seq Length - 64 64 128 128 128 256 128 128 64 512 160
a Matthew’s Correlation is reported for CoLA; F1-score for MRPC and QQP; Pearson Correlation for STS-B, and Accuracy for the rest.
b With using equation (7).
TABLE III: Results comparison between BERTbase [4], DistilBERT [46], TinyBERT [47], MobileBERT [48], PoWER-BERT [25], LayerDrop [39], FPT [34] and BERTbase with latency adjustable encoder (This Work)
Modela Method CoLA SST-2 MRPC QQP QNLI RTE MNLI-M MNLI-MM Avg Drop FLOPs Speedup
BERTbase - 52.7 93.5 88.7 71.2 91.0 68.9 84.7 83.8 0.00 22.5B 1
DistilBERT KD 51.3 91.3 87.5 70.1 89.2 59.9 82.2 - 2.74 11.3B 1.7
TinyBERT KD 51.1 93.1 87.3 71.6 90.4 70 84.6 83.2 0.40 11.3B 2.0
MobileBERTb KD 50.5 92.8 88.8 70.2 90.6 66.2 83.3 82.6 1.19 5.7B 5.5
PoWER-BERTc Pruning 52.3 92.1 88.1 70.2 90.1 67.4 83.8 83.1 0.92 - 2.9
LayerDrop Pruning - 92.5 - - 89.4 - 82.9 - 1.47 11.3 1.7
FPT Pruning - 92.6 - - 90.3 - 83.5 - 0.93 14.5 1.5
This Work Pruning 53.4 92.6 88.0 70.1 90.2 67.9 83.7 82.7 0.73 7.7B 2.9
a Matthew’s Correlation reported for CoLA; F1-score for MRPC and QQP; Pearson Correlation for STS-B; Accuracy for the rest.
b MobileBERT starting point is from BERTlarge architecture that is around three times more parameters with respect to BERTbase.
c PoWER-BERT training consists of 3 fine-tuning steps, and the model inference speedup is fixed at inference.
Table III compares the proposed method and the previously reported methods. BERTbase model ($L$=12, $h$=12, $d$=768) is used as the basis for comparison. The test results, the average accuracy drop compared to BERTbase, the average number of FLOPs, and the average inference speedup are reported for each model. As shown in the table, the proposed method achieves 2.9 times speedup with 0.73% on average accuracy drops compared to BERTbase model.
Compared to other works, DistilBERT model achieves 1.7 times speedup with around 2.74% accuracy degradation on average. In DistilBERT, a pre-trained general-purpose language presentation model must be trained from scratch with significant effort. The TinyBERT model speedup is 2.0 times with 0.4% accuracy drops on average. TinyBERT uses a two-stage learning framework that performs Transformer distillation at pre-training and fine-tuning phases with considerable excessive computational effort. The MobileBERT model speedup is 5.5 times with around 1.19% accuracy drop, but the starting point of this model is BERTlarge. The PoWER-BERT model speedup is 2.9 times with around 0.92% accuracy drop. However, the training of PoWER-BERT consists of 3 phases: ”fine-tuning, elimination configuration search, and re-training” [25] that impose extra training effort. LayerDrop and FPT speedups are 1.7 and 1.5 times, respectively, with corresponding average accuracy drops of 1.47% and 0.93%.
The table shows that the results of the proposed method are competitive and more effective than previous baseline methods across a wide range of experimental tasks. Furthermore, compared to recently reported methods, our approach is applied only during the fine-tuning phase and offers the offline-tuning property. Blank fields in the table, indicate instances where some works did not report their assessments on certain datasets.
IV-E Experimental Results of GPT-2
To assess the impact of the proposed method on the overall context of the input, the approach is implemented on the GPT-2 model following the procedures detailed in Section III-E. The training process involves utilizing 1% of the OpenWebText [82] dataset with a stride of 128 for three epochs. This methodology is employed to comprehensively evaluate how the proposed method influences the global context of the input when applied to the GPT-2 model.
Table IV presents the results of the zero-shot perplexity assessment for the proposed method in text generation tasks, offering a basis for comparison with other contemporary studies. The findings demonstrate that the proposed method effectively maintains the global context of the input sequence. The proposed Attention Context Contribution (ACC) metric performs well, ensuring that the elimination of word-vectors does not compromise the input’s global context. It is noteworthy that the original GPT-2 model, trained by OpenAI [71], benefited from more extensive hardware and dataset resources. However, our proposed architecture enables efficient adaptation with just a few small epochs. Furthermore, the proposed method exhibits superior performance when compared to the DistilGPT2 [46] and ZipGPT2 [83] models in the evaluation.
TABLE IV: Perplexity (PPL) comparison between GPT-2 [71], DistilGPT2 [46] and ZipGPT2 [83] with latency adjustable GPT-2
(This Work)
Model Wiki Text-103 LAMBADA PTB 1BWa Speedup
GPT-2 37.5 35.1 36.46 45.54 1
DistilGPT2 43.0 - - - 1.9
ZipGPT2 72.1 - - - 3.3
This Work 43.2 59.0 52.9 51.66 2.7
49.3 64.6 61.7 54.53 3.5
a The model is evaluated on 0.5% of the 1BW test set.
Table V exhibits the classification accuracy findings of the proposed method in contrast to GPT-2 and DistilGPT2. Notably, the proposed method attains an inference latency three times quicker than GPT-2, accompanied by a marginal average accuracy reduction of 0.55%. Blank fields in the tables, indicate instances where some works did not report their assessments on certain datasets.
TABLE V: Accuracy comparison between GPT-2 [71] and DistilGPT2 [46] with latency adjustable GPT-2 (This Work)
Model SST-2 IMDB Avg Drop Speedup
GPT-2 93.2 94.1 0 1
DistilGPT2 91.5 - 1.7 1.9
This Work 92.8 93.4 0.55 3.0
IV-F Experimental Results of Flan-T5
To assess the proposed method under the instruction tuning paradigm, the method is applied to the base and large versions of the Flan-T5 model which have 220M and 770M parameters respectively. The method’s performance is evaluated using different types of prompting with the MMLU and BBH benchmarks. In the evaluation procedure, no training or fine-tuning steps are required. The modified model with the original Flan-T5 model weights is studied in this section. Table VI presents the experimental results for 5-shot and 3-shot direct prompting on the MMLU and BBH benchmarks, measuring the exact-match score. As shown in the table, the model achieves up to a 3-fold encoder speedup in the MMLU benchmark with only marginal decrease in score, and in the BBH benchmark, the method functions effectively with up to a 1.5-fold encoder speedup.
TABLE VI: Experimental results of the proposed method on Flan-T5 [72] are evaluated using the MMLU [78] and BBH [79] instruction tuning datasets.
Model MMLUa BBHb
Score Encoder Speedupc Score Encoder Speedup
Flan-T5 Base (220M) $\alpha_{ER}$ = 1.00 34.05 1 30.79 1
$\alpha_{ER}$ = 0.98 34.29 1.18 30.90 1.24
$\alpha_{ER}$ = 0.95 34.13 1.41 30.37 1.46
$\alpha_{ER}$ = 0.93 34.30 1.58 28.55 1.58
$\alpha_{ER}$ = 0.90 34.13 1.83 18.92 1.83
$\alpha_{ER}$ = 0.85 33.95 2.34 - -
$\alpha_{ER}$ = 0.80 32.24 2.94 - -
Flan-T5 Large (770M) $\alpha_{ER}$ = 1.00 41.95 1 37.04 1
$\alpha_{ER}$ = 0.98 41.92 1.29 36.48 1.45
$\alpha_{ER}$ = 0.95 42.05 1.78 29.25 1.90
$\alpha_{ER}$ = 0.93 41.51 2.23 22.81 2.37
$\alpha_{ER}$ = 0.90 35.66 3.08 16.78 3.03
a MMLU includes exam questions from 57 tasks, uses 5-shot direct prompting, and measure the exact-match score.
b BBH includes 23 challenging tasks, uses 3-shot direct prompting, and measure the exact-match score.
c The proposed self-attention layer is applied to the encoder stage of the Flan-T5 model, and thus the speedup of the encoder stage is reported.
IV-G Offline-Tuning Property
After the fine-tuning phase, the trade-off between inference accuracy and speed of the model can be adjusted with the offline-tuning property. This property controls each encoder layer’s word-vector elimination rate by adjusting the Speedup-Coefficient hyper-parameter ($\alpha_{SC}$) value. Fig. 6 presents the offline-tuning property. In the following experiments, the model first fine-tuned with $\alpha_{SC}$ equal to one for a given task, and then in the offline-tuning phase, the $\alpha_{SC}$ value changes from 0.85 to 1.2 to observe its effect on inference speedup and accuracy.
Fig. 6a shows the model inference speedup for a range of $\alpha_{SC}$ hyper-parameter values on multiple tasks. The model inference speedup increased more than three times compared to BERTbase, with the $\alpha_{SC}$ value decreasing to 0.85. Fig. 6b presents the accuracy stability on multiple tasks. It shows that the inference results are kept stable over a wide range of $\alpha_{SC}$ values selections. Fig. 6c shows that the inference results are almost unchanged even if the complexity of the model reduces by 70% according to the proposed method. In these experiments, the model’s accuracy on validation sets is reported because of the GLUE test set verification limitations.
Figure 6: The proposed offline-tuning results, (a)Inference speedup compare to BERTbase versus $\alpha_{SC}$, (b)Validation set results versus $\alpha_{SC}$, (c)Validation set results versus speedup compare to BERTbase for MRPC [65], SST-2 [64], QNLI [67], IMDB [62] and MNLI-M/MM [69] tasks.
TABLE VII: Ablation study on the proposed self-attention layer with different self-attention Sort layer implementations.
Modela SST-2 MRPC MNLI-M MNLI-MM STS-B IMDB Avg
BERTbase Self-Attention 93.5 88.7 84.7 83.8 84.3 93.7 88.12
   + Eliminate-Layer 89.9 84.4 81.7 80.4 83.3 90.7 85.07
   + Eliminate-Layer + Random-Sort-Layer 84.5 80.4 63.6 64.4 55.7 87.5 72.68
   + Eliminate-Layer + SV-Sort-Layer 92.6 88.0 83.7 82.7 84.2 92.7 87.32
a F1-score for MRPC; Pearson Correlation for STS-B; Accuracy for the rest.
Two approaches can be offered to select the proper value of the $\alpha_{SC}$ hyper-parameter in the fine-tuning phase. The first approach is using equation (7) and calculating the approximate value of the $\alpha_{SC}$ based on the desired inference speedup. The second approach is to select the $\alpha_{SC}$ value using suggested Fig. 6(c) curves. After the fine-tuning phase, the speedup can be accurately adjusted with the offline-tuning property.
The stability of the proposed method’s offline perplexity (PPL) in text generation tasks is evaluated by training the model with a specific speedup coefficient and subsequently altering this value during the inference stage, measuring the model’s zero-shot perplexity. Fig. 7 illustrates the offline perplexity of the proposed method on various datasets, including WikiText-103, LAMBADA, PTB, and 1BW. The graph depicts the model’s robust stability, with a smooth perplexity curve observed across a broad range of speedup values. This underscores the efficacy of the proposed method in maintaining stable performance under varying conditions.
Figure 7: Offline Perplexity (PPL) of the proposed method on WikiText-103 [74], LAMBADA [77], PTB [75] and 1BW [76] datasets.
IV-H Ablation Studies
Table VII presents the ablation study on the proposed self-attention layer. The first row of the table presents the original BERTbase results, and the subsequent rows are the results of the self-attention layer with different types of word-vector sorting strategies.
The first approach involves removing the concluding word-vectors from the contextual sequence without the use of a sorting layer. The second approach employs the Random-Sort-Layer, which shuffles word-vectors and randomly eliminates them. The final and better strategy is the proposed strategy that incorporates the SV-Sort-Layer that arranges word-vectors according to their Score values, subsequently removing the concluding word-vectors.
As shown in the table, the first strategy shows a significant accuracy drop compared to the original BERTbase and the proposed method. The second strategy has a destructive effect on the model results, and the fine-tuning process becomes completely unstable in CoLA, QQP, QNLI, and RTE datasets not presented in the table. The third strategy is superior to other investigated strategies and proves that the proposed self-attention Sort layer is effective. In these experiments, the hyper-parameters values and the fine-tuning conditions are the same.
V Conclusion
This paper introduces a latency-adjustable Transformer encoder layer that significantly reduces the inference time of Transformer-based architectures with only a marginal decrease in accuracy. A novel Attention Context Contribution (ACC) metric is proposed to assess context contribution in each Transformer encoder layer, experimentally demonstrating that word-vector contribution diminishes in the last encoder layers. An accurate analytical speedup estimation is presented, aiding system designers in hyper-parameter selection. The powerful offline-tuning property enables adjustment of model speedup after the fine-tuning phase. The proposed method is applied to BERTbase, GPT-2, and Flan-T5 models, validated on the GLUE benchmark, the IMDB dataset, and various text generation tasks. Additionally, the method’s ability to understand language is evaluated under the instruction tuning paradigm, employing different types of prompting to assess its versatility and robustness. The suggested approach asserts that in Large Language Models (LLMs), although the complete network is vital for training, it can be truncated during the fine-tuning phase.
Acknowledgment
Thanks to Mr. Seyed Mahdi Hosseini for his assistance and comments that greatly improved our work.